# Youth as Peer Auditors: Engaging Teenagers with Algorithm Auditing of Machine Learning Applications

### v1

@@ 2.1. Youth as testers and evaluators
Several CCI studies have mentioned the importance of engaging youth in testing AI/ML models, but they often lack detailed findings on how young people evaluate models and what they can gain from the testing process (Voulgari et al., 2021; Kaspersen et al., 2022; Hitron et al., 2019; Burhans and Dantu, 2017). This is also the case in the AI/ML education literature, where training models has received the most attention (Morales-Navarro et al., 2024). The few studies that have investigated how youth test their own models show promising results. Several studies argue that when youth test models, they build hypotheses and explanations for model behaviors and also come up with new ideas for how model performance could be improved (Vartiainen et al., 2020, 2021; Druga and Ko, 2021). Recent work shows that testing can support young people in identifying issues related to data diversity, class imbalance, and data quality (Tseng et al., 2024, 2023). Yet testing is not always systematic and in-depth. Sometimes young people, after identifying cases in which models do not perform as expected, instead of trying to fix the models, change their testing behaviors (Zimmermann-Niefield et al., 2019). Other studies have shown that youth rarely test their models, only doing it when prompted by researchers, or that sometimes they think that by simply testing they can improve model performance without making changes to training datasets, model parameters, or retraining (Zimmermann-Niefield et al., 2020). A couple of studies also engage youth as testers of their peers’ models in designer-led testing activities (Dwivedi et al., 2021; Morales-Navarro et al., 2024), that is, when the designers of the applications guide their peers in the testing process. Notably, none of these studies involved youth as external evaluators that evaluate models from the outside in.
@@ 2.2. Youth as auditors of AI/ML applications
Recently, HCI researchers have started investigating how non-experts engage with algorithm auditing by involving them beyond crowdsourcing data. Studies have looked at how users engage in emergent, everyday auditing practices, without the participation of experts, on social media platforms (Shen et al., 2021). Other work has investigated approaching auditing from a socio-technical perspective by auditing both system and non-expert auditor behaviors (Lam et al., 2023, 2022). These studies highlight that non-expert audits can uncover problematic algorithmic behaviors that experts may not be able to find (Lam et al., 2022; Shen et al., 2021). Notably, all of these studies have been conducted with adults. Another important finding has been the collaborative nature of non-expert auditing which often involves sharing problematic findings with others via social media—a practice that might connect well with youth engaged in similar tasks.
@@ 2.2.3. Youth as peer auditors
To evaluate the change in youth’s auditing behaviors, before and after the workshop, we conducted a pre/post clinical interview study in which participants were presented with auditing tasks and asked to explain what they were thinking as they completed them (disessa, 2007).
@@ Appendix A Codebook
Code | Definition | Example | |||||||||||
anthropomorphizing |
“It’s thinking this is a shark” Richard | |||||||||||
auditing.breakIt |
auditing.newPerspectives |
auditing.nextSteps |
auditing.noticePatterns |
auditing.applyingKnowledge |
audtiting.placeResponsibility |
bias.age |
bias.appeareance.body |
bias.appeareance.fashion |
bias.breed |
bias.color |
bias.gender |
bias.location/context |
bias.position |
bias.race |
bias.relevancy |
bias.shape |
bias.size |
bias.socioeconomic |
dataDesign.classBalance |
dataDesign.composition |
dataDesign.context |
dataDesign.diversity |
dataDesign.edgeCase |
dataDesign.lighting |
dataDesign.quantity |
dataDesign.source |
ID.failure.explanation |
ID.failure.noExplanation |
“That’s not a shark.” Luke | |||||||||||
ID.success |
justice.exclusion |
justice.harmful |
justice.notHarmful |
justice.potentiallyHarmful |
modelDesign.features |
priorExperiences.personalBias |
priorExperiences.societalBias |
programmed |

### v3

@@ 2.1. Youth as testers and evaluators
Several CCI studies have mentioned the importance of engaging youth in testing AI/ML models, but they often lack detailed findings on how young people evaluate models and what they can gain from the testing process (Voulgari et al., 2021; Kaspersen et al., 2022; Hitron et al., 2019; Burhans and Dantu, 2017). This is also the case in the AI/ML education literature, where training models has received the most attention (Morales-Navarro et al., 2024). The few studies that have investigated how youth test their own models show promising results. Several studies argue that when youth test models, they build hypotheses and explanations for model behaviors and also come up with new ideas for how model performance could be improved (Vartiainen et al., 2020, 2021; Druga and Ko, 2021). Recent work shows that testing can support young people in identifying issues related to data diversity, class imbalance, and data quality (Tseng et al., 2024, 2023). Yet testing is not always systematic and in-depth. Sometimes young people, after identifying cases in which models do not perform as expected, instead of trying to fix the models, change their testing practices (Zimmermann-Niefield et al., 2019). Other studies have shown that youth rarely test their models, only doing it when prompted by researchers, or that sometimes they think that by simply testing they can improve model performance without making changes to training datasets, model parameters, or retraining (Zimmermann-Niefield et al., 2020). A couple of studies also engage youth as testers of their peers’ models in designer-led testing activities (Dwivedi et al., 2021; Morales-Navarro et al., 2024), that is, when the designers of the applications guide their peers in the testing process. Notably, none of these studies involved youth as external evaluators that evaluate models from the outside in.
@@ 2.2. Youth as auditors of AI/ML applications
Recently, HCI researchers have started investigating how non-experts engage with algorithm auditing by involving them beyond crowdsourcing data. Studies have looked at how users engage in emergent, everyday auditing practices, without the participation of experts, on social media platforms (Shen et al., 2021). Other work has investigated approaching auditing from a socio-technical perspective by auditing both system and non-expert auditor practices (Lam et al., 2023, 2022). These studies highlight that non-expert audits can uncover problematic algorithmic behaviors that experts may not be able to find (Lam et al., 2022; Shen et al., 2021). Notably, all of these studies have been conducted with adults. Another important finding has been the collaborative nature of non-expert auditing which often involves sharing problematic findings with others via social media—a practice that might connect well with youth engaged in similar tasks.
@@ 2.2.3. Youth as peer auditors
To evaluate the potential benefit of peer auditing activities, before and after the workshop, we conducted a pre/post clinical interview study in which participants were presented with auditing tasks and asked to explain what they were thinking as they completed them (disessa, 2007).
@@ Appendix A Codebook
anthropomorphizing
Definition: Attaching human characteristics to the model.
Example: “It’s thinking this is a shark” Richard
auditing.breakIt
Definition: Indicating that auditing involves finding moments when the project breaks.
Example: “Challenging it [the project] to see what would break it ” Iván
auditing.newPerspectives
Definition: Voicing that auditors provide new perspectives on how the project works.
Example: “it is helpful to get feedback and other perspectives from other people. Who may see things that we have not seen.” Luke
auditing.nextSteps
Definition: Coming up with next steps to improve model performance.
Example: “Add more variety in the data. Like these are all the same kind of rabbit.” Kayla
auditing.noticePatterns
Definition: Identifying patterns in the outputs to build explanations for model behaviors.
Example: “which I feel like that’s kind of customary for DALL-E because I feel like that the images it’s been fed with are more probably just white people instead of diverse like images.” Iván
auditing.applyingKnowledge
Definition: Indicating that insights gained while auditing peers’ projects can be helpful to improve their own projects.
Example: “I felt like just going in, and being able to see other people’s projects and see what they can improve on. It’s like, what can I improve on it is like, your telling somebody else, what they can improve on and you can turn around and improve that yourself.” Jerome
audtiting.placeResponsibility
Definition: Referring to something or someone as responsible for the outputs.
Example: “I think it shows more of like human bias than like, AI bias maybe that that’s like what like the like, because if it’s like trained off of like pictures, like you were showing, and that’s kind of like the pictures that it’s been like that it’s seen being put up on the internet.” Jackie Star
bias.age
Definition: Identifying potential age related biases in model outputs.
Example: “it’s all like let’s say young woman like for like the other ones there’s like more young people but like they should I feel like they should I add like more older people are like mid-age these are these people look really young” Walter
bias.appearance.body
Definition: Identifying potential body appearance related biases in model outputs.
Example: “it’s all like let’s say young woman like for like the other ones there’s like more young people but like they should I feel like they should I add like more older people are like mid-age these are these people look really young” Walter
bias.appearance.body
Definition: Identifying potential fashion related biases in model outputs.
Example: “like the general like, in the lab, mixing stuff together, lab coats, goggles gloves. Again, is mainly just just white people.” Jerome
bias.breed
Definition: Identifying potential breed related biases in model outputs.
Example: “Yeah like the Corgi you can see is all… the Corgi is [identified as] a dog and the German Shepherd as the cat. Why would it think that?” Kayla
bias.color
Definition: Identifying potential breed related biases in model outputs.
Example: “Yeah like the Corgi you can see is all… the Corgi is [identified as] a dog and the German Shepherd as the cat. Why would it think that?” Kayla
bias.gender
Definition: Identifying potential gender related biases in model outputs.
Example: “The models were all white girls with straight hair..” Jackie Star
bias.location/context
Definition: Identifying potential location/context related biases in model outputs.
Example: “All the dolphins are jumping on the water none of this whales are so it’s creating a bias towards the dolphins because they’re all technically in the same act that having the same like actions and as dolphins.” Fatimah
bias.position
Definition: Identifying potential position related biases in model outputs.
Example: “These cats are mainly perched up. Yeah, I feel like in the same position.” Emily
bias.race
Definition: Identifying potential race related biases in model outputs.
Example: “For the doctor that was always white male, or any other professions like lawyer teacher well teachers mean female or like librarians female but it was majority white people and no black or any race.” Fatimah
bias.relevancy
Definition: Identifying potential relevancy related biases in model outputs.
Example: “This was probably how what a good student was like defined as back in the day.” Stephanie
bias.shape
Definition: Identifying potential shape related biases in model outputs.
Example: “Because I had just like a prominent shape like yeah, shape or triangles and cones. It’s easy to see compared to the other two, which can be kind of a bit similar. It made the [blackberry] look like strawberries.” Kayla
bias.size
Definition: Identifying potential size related biases in model outputs.
Example: “this thing just thinks everything that has big ears as a cat.” Richard
bias.socioeconomic
Definition: Identifying potential biases related to socioeconomic status in model outputs.
Example: “They’re all like white again, like probably middle class… upper middle class white people… mainly white men or boys I guess. ” Kayla
dataDesign.classBalance
Definition: Inferring data composition related issues in the design of training datasets.
Example: “I would add more pictures of like sharks. Like zoomed out like the whole body color, good lighting and the whale add more like out of water where it is jumping maybe like the dolphins nothing because it seems to get like a dolphins pretty accurately.” Walter
dataDesign.context
Definition: Inferring data context related issues in the design of training datasets.
Example: “There may be like hands involved or like, like a background, like a lot of greenery so it knows what’s the difference… because this one’s just covered, the background is just white. So there’s nothing, nothing in the background.” Walter
dataDesign.diversity
Definition: Inferring data diversity related issues in the design of training datasets.
Example: “More representation definitely just different environments different people different skin colors, races genders and even for this maybe like same sex couples that could be something” Fatimah
dataDesign.edgeCase
Definition: Identifying potential edge cases.
Example: “I don’t think it’s really used to like pencils or markers that are like black and white because all the other examples are brightly colored yeah.” Jackie Star
dataDesign.lighting
Definition: Inferring lighting related issues in the design of training datasets.
Example: “Like better lighting for this one, because they’re all like underwater.” Kayla
dataDesign.quantity
Definition: Inferring data quantity related issues in the design of training datasets.
Example: “We would add more hands to holding new markers, we will do some darker markers. There’s only like two darker markers.” Emily
dataDesign.source
Definition: Inferring issues related to the sources of data used in training datasets.
Example: “Maybe based off of like stock images, like that’s what it was trained on.” Kayla
ID.failure.explanation
Definition: Identifying cases in which the model did not perform as expected and providing an explanation for such performance.
Example: “Yeah, it’s so they go based off of the sort of since the cat is white. It goes to the bunny since all of the pictures of the bunny are white.” Luke
ID.failure.noExplanation
Definition: Identifying cases in which the model did not perform as expected without providing an explanation for such performance.
Example: “That’s not a shark.” Luke
ID.success
Definition: Identifying cases in which the model performs as expected.
Example: “They all have like the big ears. I think that’s why they got these two right because they have big dog ears.” Kayla
justice.exclusion
Definition: Arguing that system outputs exclude some people.
Example: “Because people can look at this, these pictures of doctors and rich people and see that they look nothing like them? And then feel discouraged? As if Oh, that’s not an opportunity that I can convey?” Luke
justice.harmful
Definition: Arguing that system outputs produce harm.
Example: “Mainly the librarian one is harmful because it shows a bunch of women… it’s not like men can be librarians, ”don’t do that… it’s not a masculine job, you shouldn’t have that job if you’re a man.” I guess it’s kind of saying.” Kayla
justice.notHarmful
Definition: Arguing that system outputs do not produce harm.
Example: “I think if you’re getting harmed by an AI, I don’t know. That’s more of a personal problem.” Richard
justice.potentiallyHarmful
Definition: Arguing that system outputs could be potentially harmful.
Example: “I feel like at this point right now, it’s not harmful but like, as it evolves, it will be is like beautiful. I feel like no one’s actually going to compare themselves to these awful images, and be like, Wow, if I don’t if I don’t look like this person, my God and of the world, but I feel like as it evolves, it will be more harmful if these issues aren’t addressed by adding more diversity.“ Iván
modelDesign.features
Definition: Inferring parameters or features used by the model.
Example: “I don’t think it’s taking in color, it doesn’t care if the strawberry is white or whatever the color of strawberries it’s more looking at the texture.” Richard
priorExperiences.personalBias
Definition: Considering personal biases in evaluating system behaviors.
Example: “It makes sense that it will all be woman. I’ve personally never heard of a male librarian… I really haven’t.” Kayla
priorExperiences.societalBias
Definition: Considering societal biases in evaluating system behaviors.
Example: “What about like cook because I feel like that could go in either direction. Yeah, because it could be stereotyped with women being in the kitchen…” Iván
programmed
Definition: Indicating that the machine has been programmed to work in a certain way.
Example: “mean, I’ve seen… I saw like this one thing, I don’t remember where it was, it was like, it asked the chat GPT to do a prompt for a smart scientist. And it just put out that the scientists had to be white. And I found that was kind of interesting, because it was that was the biases that it was programmed with, when that’s not true.” Iván